OpenStudy (butterflydreamer):

If dy/dx = 1+y and when x =0 , y=2, find y as a function of x

OpenStudy (irishboy123):

separate the variables, ie x & y, and integrate.

OpenStudy (chaise):

\[\frac{dy}{dx}=1+y\] \[\frac{dy}{1+y}=dx\] \[\frac{1}{1+y}dy=dx\] Intergrate both sides and solve, don't forget your integration constant c, subsitute in your values for x and y, solve for c and rearrange for y in terms for x. Do you need more help?

OpenStudy (butterflydreamer):

aahhhh i see! Completely forgot that you could do that. Thank you @chaise and @IrishBoy123 :) !!

OpenStudy (chaise):

No problem, do you know how to integrate: \[\frac{1}{1+y} dy\]

OpenStudy (butterflydreamer):

yeah that's just log (y+1) + c right?